OpenStudy (anonymous):

Solve the exponential equation. 9^8x = 27 I don't understand how to solve this. If you know I will give medal (: A) 5/8 B) 3/16 C) 9 D) 16/3

OpenStudy (cwrw238):

hint take logarithns

OpenStudy (mathmale):

Hint: rewrite both the 9 and the 27 as powers of 3:\[(3^{2})^{8x}=3^{3}\] since y ou now have the same base on both sides, it's easy to solve for x. Try it. If the resulting x doesn't match any of your answers, double check to ascertain that you've copied the problem down correctly.

OpenStudy (anonymous):

so the logarithm would look like this right? log8^27 = x

OpenStudy (cwrw238):

or you can convert as follows - i was just going to do what mathmale has just written

OpenStudy (anonymous):

alright. I will solve and then tell you what I got. Then you or mathmale will tell me if I'm right.

OpenStudy (cwrw238):

no 8x *log 9 = log 27

OpenStudy (mathmale):

If you, usernae, choose to use logs to solve this problem for x, you'd end up with 8x log 9 = log 27. Solve for x.

OpenStudy (anonymous):

so do I divide log 9 from log 27 in order to be left with x. Then simplify from there?

OpenStudy (mathmale):

Yes, that's an appropriate approach. Then you'll need to isolate x.

OpenStudy (anonymous):

my answer is x = 3/8

OpenStudy (mathmale):

Please check that result by substitution.

OpenStudy (cwrw238):

no thats not the answer show us your working

OpenStudy (anonymous):

I divided log27/log9

OpenStudy (cwrw238):

x should be = [log 27 / log 9] / 8

OpenStudy (anonymous):

then I isolated the x by dividing both sides by 8

OpenStudy (cwrw238):

right so far

OpenStudy (cwrw238):

yep

OpenStudy (anonymous):

but that's not one of the answer choices. However, when I plugged in the logarithm into wolfram it gave me 3/16. Where did the 16 come from? Do you know?

OpenStudy (cwrw238):

what did you get log 27 / log 9?

OpenStudy (anonymous):

yes I did (:

OpenStudy (mathmale):

If \[9^{8x}=27,\] \[(3^{2})^{8x}=3^{3}\] is also true. Then 16x=3, and x = ???

OpenStudy (mathmale):

Logs should also lead to the correct answer.

OpenStudy (cwrw238):

log27/lgo9 = 3/2 divide this by 8 gives 3/2 * 1/8 = 3/16

OpenStudy (mathmale):

I advocated this particular solution some time ago, but read it incorrectly and tried to solve 6x=3 for x, whereas it should be 16x=3.

OpenStudy (anonymous):

but wait. 27/9 = 3.

OpenStudy (mathmale):

While 27/9=3, you are actually dividing (log 27) by (log 9).

OpenStudy (cwrw238):

you need to divide the logs not the numbers log 27 / log 9 = 1.5 or 3/2

OpenStudy (mathmale):

And that 's not 3.

OpenStudy (anonymous):

ohhh okay then. I get it now. so the answer is 3/16. Thank you @cwrw238! and thank you mathmale! Listen if it's not too much, I have one more question.

OpenStudy (mathmale):

Post it separately, please!